# RAG-R1: Incentivizing the Search and Reasoning Capabilities of LLMs through Multi-query Parallelism

### v5

@@ 1 Introduction
Large Language Models (LLMs) (Survey_of_LLMs; deepseek-r1) have demonstrated remarkable capabilities in various domains, including mathematical reasoning, question answering, and code generation. However, the knowledge encoded in these models is static, limiting their adaptability. As a result, LLMs are susceptible to producing hallucinated or outdated responses (Survey_of_Hallucination; RAG_Reduce_Hallucination; self-rag) when dealing with complex or real-time issues, compromising their reliability. Therefore, it is essential to equip LLMs with access to external knowledge to ensure more accurate and grounded responses.
@@
Retrieval-Augmented Generation (RAG) (RAG_A_Survey; RAG_Meeting_LLMs) is a widely adopted approach to address this issue, broadening the model’s capability boundaries by integrating external knowledge into the generation process. Early efforts (ITER-RETGEN; In_Context_RAG) in this area concentrated on prompt-based strategies that guide LLMs through question decomposition, query rewriting and multi-turn retrieval. While effective, these approaches are constrained by the limitations inherent in prompt engineering. Recent advances (IRCoT; Search-o1; self-rag; DeepRAG; CoRAG) have increasingly emphasized the integration of reasoning capabilities within RAG. These approaches combine RAG with Chain of Thought (CoT) for step-by-step retrieval, or automatically generate intermediate retrieval chains that incorporate reasoning and employ Supervised Fine-Tuning (SFT) to enable learning of retrieval and reasoning. However, new insights (SFT_Memorizes) indicate that these methods may cause models to memorize solution paths, thus constraining their generalization to novel scenarios.
@@
Recently, reinforcement learning (RL) (PPO) has demonstrated great potential in improving LLM performance by enhancing the reasoning capability. Reasoning models such as OpenAI-o1 (openai-o1) and Deepseek-R1 (deepseek-r1) indicate that utilizing RL with outcome-based rewards can enhance the model’s performance in mathematical and logical reasoning tasks. Within this paradigm, several works have explored enhancing the model’s search ability during reasoning through RL. R1-Searcher (R1-Searcher) proposes a novel two-stage outcome-based RL approach designed to enhance the search capability of LLMs. Search-R1 (Search-R1) introduces a novel RL framework that enables LLMs to interact with search engines in an interleaved manner with their own reasoning. Despite significant improvements, these RL-based methods struggle with stable training due to the restricted capabilities of cold-start models (deepseek-r1). Furthermore, existing methods generate only a single search query whenever external retrieval is required, which presents two significant challenges: (1) Prohibitive Latency from Serial Execution. The single-query architecture mandates a serial execution model for multi-hop reasoning, where each step must await the completion of the preceding retrieval-inference cycle. Consequently, latency accumulates with each turn, rendering the model impractical for latency-sensitive, real-world applications. (2) Inherent Brittleness and Knowledge Confinement. The single-query approach is inherently brittle, as its entire reasoning trajectory depends on the success of each sequential step. An early, suboptimal query can lock the model into an unrecoverable path—a critical failure mode. This single-threaded process confines the model to a narrow slice of external knowledge, leaving it unable to navigate user ambiguity, explore alternative reasoning paths, or recover from an initial misstep. This fragility ultimately compromises its reasoning robustness and final performance. We conducted a straightforward experiment based on Qwen2.5-72B-Instruct (qwen2.5) following Search-R1 to validate the aforementioned challenges. As illustrated in Figure 1, we evaluated the model’s performance and average retrieval count on two Multi-Hop Question Answering benchmarks: HotpotQA (yang2018hotpotqa) and 2WikiMultiHopQA (xanh2020_2wikimultihop), comparing the generation of a single query to multiple queries in scenarios necessitating retrieval. The findings indicate that the multi-query method enhances the model’s performance and decreases retrieval iterations compared to the single-query approach, highlighting the limitations inherent in the single-query mode.
@@
We conduct extensive experiments based on Qwen2.5-7B-Instruct (qwen2.5) to verify the effectiveness of our proposed method. The results demonstrate its effectiveness, which achieves state-of-the-art performance on seven question-answering benchmarks. In particular, our method utilizing multiple-query parallelism outperforms the strongest baseline by up to 13.7% and decreases inference time by 11.1%. The contributions of this work are summarized as follows:
@@ 2 Related Work
Retrieval-Augmented Generation RAG enhances LLMs outputs by integrating external information with internal knowledge. Initial RAG methods employed static pipelines, coupling retrieved documents with LLMs via prompt engineering (replug; ralm; query-decomposing; multi-aspect-query) to control generation. While effective for information-sourcing tasks, these approaches struggle with complex, multi-faceted queries requiring iterative retrieval. Subsequent research developed dynamic RAG frameworks (adaptive-rag) that utilize LLM feedback for adaptive retrieval and generation control. Complementary efforts have focused on improving knowledge representation of retrieved documents (graph-rag; light-rag) from augmentation perspectives. Most recently, advances (Search-R1; Search-o1; R1-Searcher; zerosearch; masksearch) demonstrate significant scalability potential by training interaction trajectories to optimize LLM’s use of search engines for complex problem-solving queries. Following these advances, our method further enhances the interaction between LLMs and retrievers by leveraging LLM’s reasoning abilities.
Learning to Search Recent advances in Learning to Search have evolved from static query and generation template via SFT to dynamic and reward-driven search using RL. SFT-based approaches have demonstrated success in improving LLM capabilities in instruction following (vif-rag; self-play; retrollm), robustness (dpa-rag), and domain-specific adaptation (raft). However, these methods are limited by a tendency to memorize solution paths, which constrains generalization and scalability. RL-based methods (openai-o1; deepseek-r1; grpo; qwen3) offer a promising alternative by enabling autonomous reasoning and decision-making. Recent works (Search-o1; R1-Searcher; Search-R1) further integrate LLMs into real-time search workflows, allowing interaction with live search engines. Although these methods support autonomous search engine invocation, the efficacy of search-as-a-tool remains underexplored. To address this, our approach introduces a two-stage training framework that enables parallel query generation, with the aim of retrieving more comprehensive and diverse document contexts.
@@ 3.1 Format Learning Supervised Fine-Tuning
To equip LLMs with the ability to adaptively leverage internal and external knowledge during reasoning and respond in a think-then-search format, we first generate samples that integrate reasoning and search. Specifically, the system instruction guides the LLM to perform reasoning between <<think>> and <</think>> whenever it receives new information. After completing the reasoning process, the LLM can generate search query between the designated search call tokens, <<search>> and <</search>>, whenever external retrieval is necessary. The system will subsequently extract the search query and request an external search engine to retrieve relevant documents. The retrieved information is then appended to the existing sequence, enclosed within special retrieval tokens, <<information>> and <</information>>, providing additional context for the next generation step. This process continues iteratively until reaching the maximum number of retrieval or the model generates a final answer enclosed within the special answer tokens, <<answer>> and <</answer>>. The system instruction follows Search-R1 and is shown in Table 1.
@@
We use Qwen2.5-72B-Instruct as the generation model and employ a portion of the HotpotQA (yang2018hotpotqa) training dataset for generation tasks.
@@ 3.2 Retrieval-Augmented Reinforcement Learning
where $\pi_{\theta}$ and $\pi_{ref}$ represent the policy model and the reference model, respectively, both of which are initialized from the SFT model, $r_{\phi}$ is the reward function and $\mathbb{D}_{KL}$ is KL-divergence. $x$ denote input samples drawn from the dataset $D$, and $y$ represent the generated outputs which is sampled from the policy model $\pi_{\theta}$ and retrieved from the retrieval system $\mathcal{R}$. The rollout process follows the same procedure detailed in Section 3.1. We employ the Proximal Policy Optimization (PPO) (PPO) algorithm to optimize the policy model, which is widely used in reinforcement learning due to its efficiency and reliability.
@@
Reward Modeling. The reward function serves as the primary training signal, which decides the optimization direction of RL. Inspired by deepseek-r1, we adopt a rule-based reward system that includes final answer rewards to assess the accuracy of responses. For instance, in factual reasoning tasks, we adopt rule-based criteria such as exact string matching (EM) to evaluate correctness:
@@
where $a_{pre}$ represents the extracted final answer from response $y$ and $a_{gold}$ denotes the ground truth answer. This method focuses on ensuring that the final answer matches the expected truth precisely. We avoid incorporating format rewards because the SFT model has already learned to respond in a think-then-search format. This strategy enables us to concentrate on the factual accuracy of the responses, leveraging the inherent structured approach of the model without imposing additional format-based conditions. Furthermore, following deepseek-r1, we do not apply neural reward models to avoid reward hacking and additional computational cost.
@@ 5.1 Experimental Settings
Datasets and Evaluation Metrics. We evaluate our models on seven benchmark datasets: (1) General Question Answering: NQ (Natural_Questions), TriviaQA (JoshiTriviaQA2017), and PopQA (mallen2023trustlanguagemodelsinvestigating). (2) Multi-Hop Question Answering: HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (xanh2020_2wikimultihop), Musique (trivedi2021musique), and Bamboogle (press-etal-2023-measuring). HotpotQA is in-domain benchmark since a portion of its training sets is used for training while others serve as out-of-domain benchmarks to assess our model’s generalization capabilities. For evaluation metric, we utilize Exact Match (EM) score following rankrag and Retrieval Count (RC).
Baselines. We compare our method against the following baselines: (1) Direct Inference: Generate answers based on the model’s internal knowledge. (2) Standard RAG: Traditional retrieval-augmented generation systems. (3)RAG-CoT Methods: Integration of retrieval-augmented generation with chain-of-thought prompts, such as IRCoT (IRCoT) and Search-o1 (Search-o1). (4) RAG-RL Methods: Utilizing reinforcement learning to allow the model to autonomously perform retrieval during inference, such as Search-R1 (Search-R1) and R1-Searcher (R1-Searcher).
Implementation Details. We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (KILT) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (bge_m3) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison. For evaluation, the number of retrieval iterations was left uncapped.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (GAE) with parameters $\lambda=1$ and $\gamma=1$. All the trainings are conducted on a single node with 8 A100 GPUs.
@@ Appendix A Implementation Details
We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (KILT) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (bge_m3) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (GAE) with parameters $\lambda=1$ and $\gamma=1$. All the trainings are conducted on a single node with 8 A100 GPUs. We configure the batch settings as follows: a total batch size of 512, a mini-batch size of 128, and a micro-batch size of 32. The maximum sequence length is 8,192 tokens, with a maximum response length of 500 and retrieved content lengths set to a maximum of 1000 tokens for each search query.

### v6

@@ 1 Introduction
Large Language Models (LLMs) (Zhao et al. 2023; Guo et al. 2025) have demonstrated remarkable capabilities in various domains, including mathematical reasoning, question answering, and code generation. However, the knowledge encoded in these models is static, limiting their adaptability. As a result, LLMs are susceptible to producing hallucinated or outdated responses (Ji et al. 2023; Shuster et al. 2021; Asai et al. 2024) when dealing with complex or real-time issues, compromising their reliability. Therefore, it is essential to equip LLMs with access to external knowledge to ensure more accurate and grounded responses.
@@
Retrieval-Augmented Generation (RAG) (Gao et al. 2023; Fan et al. 2024) is a widely adopted approach to address this issue, broadening the model’s capability boundaries by integrating external knowledge into the generation process. Early efforts (Shao et al. 2023; Ram et al. 2023a) in this area concentrated on prompt-based strategies that guide LLMs through question decomposition, query rewriting and multi-turn retrieval. While effective, these approaches are constrained by the limitations inherent in prompt engineering. Recent advances (Trivedi et al. 2022a; Li et al. 2025; Asai et al. 2024; Guan et al. 2025; Wang et al. 2025) have increasingly emphasized the integration of reasoning capabilities within RAG. These approaches combine RAG with Chain of Thought (CoT) for step-by-step retrieval, or automatically generate intermediate retrieval chains that incorporate reasoning and employ Supervised Fine-Tuning (SFT) to enable learning of retrieval and reasoning. However, new insights (Chu et al. 2025) indicate that these methods may cause models to memorize solution paths, thus constraining their generalization to novel scenarios.
@@
Recently, reinforcement learning (RL) (Schulman et al. 2017) has demonstrated great potential in improving LLM performance by enhancing the reasoning capability. Reasoning models such as OpenAI-o1 (Jaech et al. 2024) and Deepseek-R1 (Guo et al. 2025) indicate that utilizing RL with outcome-based rewards can enhance the model’s performance in mathematical and logical reasoning tasks. Within this paradigm, several works have explored enhancing the model’s search ability during reasoning through RL. R1-Searcher (Song et al. 2025) proposes a novel two-stage outcome-based RL approach designed to enhance the search capability of LLMs. Search-R1 (Jin et al. 2025) introduces a novel RL framework that enables LLMs to interact with search engines in an interleaved manner with their own reasoning. Despite significant improvements, these RL-based methods struggle with stable training due to the restricted capabilities of cold-start models (Guo et al. 2025). Furthermore, existing methods generate only a single search query whenever external retrieval is required, which presents two significant challenges: (1) Prohibitive Latency from Serial Execution. The single-query architecture mandates a serial execution model for multi-hop reasoning, where each step must await the completion of the preceding retrieval-inference cycle. Consequently, latency accumulates with each turn, rendering the model impractical for latency-sensitive, real-world applications. (2) Inherent Brittleness and Knowledge Confinement. The single-query approach is inherently brittle, as its entire reasoning trajectory depends on the success of each sequential step. An early, suboptimal query can lock the model into an unrecoverable path—a critical failure mode. This single-threaded process confines the model to a narrow slice of external knowledge, leaving it unable to navigate user ambiguity, explore alternative reasoning paths, or recover from an initial misstep. This fragility ultimately compromises its reasoning robustness and final performance. We conducted a straightforward experiment based on Qwen2.5-72B-Instruct (Team 2024) following Jin et al. (2025) to validate the aforementioned challenges. As illustrated in Figure 1, we evaluated the model’s performance and average retrieval count on two Multi-Hop Question Answering benchmarks: HotpotQA (Yang et al. 2018) and 2WikiMultiHopQA (Ho et al. 2020), comparing the generation of a single query to multiple queries in scenarios necessitating retrieval. The findings indicate that the multi-query method enhances the model’s performance and decreases retrieval iterations compared to the single-query approach, highlighting the limitations inherent in the single-query mode.
@@
We conduct extensive experiments based on Qwen2.5-7B-Instruct (Team 2024) to verify the effectiveness of our proposed method. The results demonstrate its effectiveness, which achieves state-of-the-art performance on seven question-answering benchmarks. In particular, our method utilizing multiple-query parallelism outperforms the strongest baseline by up to 13.7% and decreases inference time by 11.1%. The contributions of this work are summarized as follows:
@@ 2 Related Work
Retrieval-Augmented Generation RAG enhances LLMs outputs by integrating external information with internal knowledge. Initial RAG methods employed static pipelines, coupling retrieved documents with LLMs via prompt engineering (Shi et al. 2023; Ram et al. 2023b; Lin et al. 2023; Korikov et al. 2024) to control generation. While effective for information-sourcing tasks, these approaches struggle with complex, multi-faceted queries requiring iterative retrieval. Subsequent research developed dynamic RAG frameworks (Jeong et al. 2024) that utilize LLM feedback for adaptive retrieval and generation control. Complementary efforts have focused on improving knowledge representation of retrieved documents (Edge et al. 2024; Guo et al. 2024) from augmentation perspectives. Most recently, advances (Jin et al. 2025; Li et al. 2025; Song et al. 2025; Sun et al. 2025; Wu et al. 2025) demonstrate significant scalability potential by training interaction trajectories to optimize LLM’s use of search engines for complex problem-solving queries. Following these advances, our method further enhances the interaction between LLMs and retrievers by leveraging LLM’s reasoning abilities.
Learning to Search Recent advances in Learning to Search have evolved from static query and generation template via SFT to dynamic and reward-driven search using RL. SFT-based approaches have demonstrated success in improving LLM capabilities in instruction following (Dong et al. 2024b, a; Li et al. 2024), robustness (Dong et al. 2025), and domain-specific adaptation (Zhang et al. 2024). However, these methods are limited by a tendency to memorize solution paths, which constrains generalization and scalability. RL-based methods (Jaech et al. 2024; Guo et al. 2025; Shao et al. 2024; Yang et al. 2025) offer a promising alternative by enabling autonomous reasoning and decision-making. Recent works (Li et al. 2025; Song et al. 2025; Jin et al. 2025) further integrate LLMs into real-time search workflows, allowing interaction with live search engines. Although these methods support autonomous search engine invocation, the efficacy of search-as-a-tool remains underexplored. To address this, our approach introduces a two-stage training framework that enables parallel query generation, with the aim of retrieving more comprehensive and diverse document contexts.
@@ 3.1 Format Learning Supervised Fine-Tuning
To equip LLMs with the ability to adaptively leverage internal and external knowledge during reasoning and respond in a think-then-search format, we first generate samples that integrate reasoning and search. Specifically, the system instruction guides the LLM to perform reasoning between <<think>> and <</think>> whenever it receives new information. After completing the reasoning process, the LLM can generate search query between the designated search call tokens, <<search>> and <</search>>, whenever external retrieval is necessary. The system will subsequently extract the search query and request an external search engine to retrieve relevant documents. The retrieved information is then appended to the existing sequence, enclosed within special retrieval tokens, <<information>> and <</information>>, providing additional context for the next generation step. This process continues iteratively until reaching the maximum number of retrieval or the model generates a final answer enclosed within the special answer tokens, <<answer>> and <</answer>>. The system instruction follows Jin et al. (2025) and is shown in Table 1.
@@
We use Qwen2.5-72B-Instruct as the generation model and employ a portion of the HotpotQA (Yang et al. 2018) training dataset for generation tasks.
@@ 3.2 Retrieval-Augmented Reinforcement Learning
where $\pi_{\theta}$ and $\pi_{ref}$ represent the policy model and the reference model, respectively, both of which are initialized from the SFT model, $r_{\phi}$ is the reward function and $\mathbb{D}_{KL}$ is KL-divergence. $x$ denote input samples drawn from the dataset $D$, and $y$ represent the generated outputs which is sampled from the policy model $\pi_{\theta}$ and retrieved from the retrieval system $\mathcal{R}$. The rollout process follows the same procedure detailed in Section 3.1. We employ the Proximal Policy Optimization (PPO) (Schulman et al. 2017) algorithm to optimize the policy model, which is widely used in reinforcement learning due to its efficiency and reliability.
@@
Reward Modeling. The reward function serves as the primary training signal, which decides the optimization direction of RL. Inspired by Guo et al. (2025), we adopt a rule-based reward system that includes final answer rewards to assess the accuracy of responses. For instance, in factual reasoning tasks, we adopt rule-based criteria such as exact string matching (EM) to evaluate correctness:
@@
where $a_{pre}$ represents the extracted final answer from response $y$ and $a_{gold}$ denotes the ground truth answer. This method focuses on ensuring that the final answer matches the expected truth precisely. We avoid incorporating format rewards because the SFT model has already learned to respond in a think-then-search format. This strategy enables us to concentrate on the factual accuracy of the responses, leveraging the inherent structured approach of the model without imposing additional format-based conditions. Furthermore, following Guo et al. (2025), we do not apply neural reward models to avoid reward hacking and additional computational cost.
@@ 5.1 Experimental Settings
Datasets and Evaluation Metrics. We evaluate our models on seven benchmark datasets: (1) General Question Answering: NQ (Kwiatkowski et al. 2019), TriviaQA (Joshi et al. 2017), and PopQA (Mallen et al. 2022). (2) Multi-Hop Question Answering: HotpotQA (Yang et al. 2018), 2WikiMultiHopQA (Ho et al. 2020), Musique (Trivedi et al. 2022b), and Bamboogle (Press et al. 2022). HotpotQA is in-domain benchmark since a portion of its training sets is used for training while others serve as out-of-domain benchmarks to assess our model’s generalization capabilities. For evaluation metric, we utilize Exact Match (EM) score following Yu et al. (2024) and Retrieval Count (RC).
Baselines. We compare our method against the following baselines: (1) Direct Inference: Generate answers based on the model’s internal knowledge. (2) Standard RAG: Traditional retrieval-augmented generation systems. (3)RAG-CoT Methods: Integration of retrieval-augmented generation with chain-of-thought prompts, such as IRCoT (Trivedi et al. 2022a) and Search-o1 (Li et al. 2025). (4) RAG-RL Methods: Utilizing reinforcement learning to allow the model to autonomously perform retrieval during inference, such as Search-R1 (Jin et al. 2025) and R1-Searcher (Song et al. 2025).
Implementation Details. We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (Petroni et al. 2020) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (Chen et al. 2024) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison. For evaluation, the number of retrieval iterations was left uncapped.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (Schulman et al. 2015) with parameters $\lambda=1$ and $\gamma=1$. All the trainings are conducted on a single node with 8 A100 GPUs.
@@ Appendix A Implementation Details
We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (Petroni et al. 2020) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (Chen et al. 2024) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (Schulman et al. 2015) with parameters $\lambda=1$ and $\gamma=1$. All the trainings are conducted on a single node with 8 A100 GPUs. We configure the batch settings as follows: a total batch size of 512, a mini-batch size of 128, and a micro-batch size of 32. The maximum sequence length is 8,192 tokens, with a maximum response length of 500 and retrieved content lengths set to a maximum of 1000 tokens for each search query.